(ii) Find the equation of the straight line perpedicular to line
y =x/2 + 1,and passing through the point (4,5).

Answer by Alan3354(69443)

(ii) Find the equation of the straight line perpedicular to line
y =x/2 + 1,and passing through the point (4,5).
----------
The slope, m, of lines perpendicular is the negative inverse fo the given line.
-------------
Then, y-y1 = m*(x-x1) where (x1,y1) is the point.